OpenStudy (anonymous):

Solve: 6x^2 - 12x = 0

OpenStudy (sgadi):

\[x=0\] or \[x=2\]

OpenStudy (anonymous):

To illustrated @sgadi's solution: 6(x*x - 2x) = 0 x*x - 2x = 0 x(x-2) = 0 So x = 0/(x-2) x = 0 Or x-2 = 0/x x-2 = 0 x = 2